### v1

# How Grounded is Wikipedia? A Study on Structured Evidential Support

Abstract
Wikipedia is a critical resource for modern NLP, serving as a rich repository of up-to-date and citation-backed information on a wide variety of subjects. The reliability of Wikipedia—its groundedness in its cited sources—is vital to this purpose. This work provides a quantitative analysis of the extent to which Wikipedia is so grounded and of how readily grounding evidence may be retrieved. To this end, we introduce PeopleProfiles  [^1] https://github.com/wgantt/people-profiles—a large-scale, multi-level dataset of claim support annotations on Wikipedia articles of notable people. We show that $\sim 20\%$ of claims in Wikipedia lead sections are unsupported by the article body; $\sim 27\%$ of annotated claims in the article body are unsupported by their (publicly accessible) cited sources; and $>80\%$ of lead claims cannot be traced to these sources via annotated body evidence. Further, we show that recovery of complex grounding evidence for claims that are supported remains a challenge for standard retrieval methods.
1 Introduction
Long an essential ingredient for LLM pretraining, Wikipedia is now widely used during inference as a repository of high-quality, citation-backed information for RAG applications (Lewis et al., 2020; Chen et al., 2020; Fan et al., 2024, i.a.). In parallel, Wikipedia has played a major role in advancing automated fact or claim verification (Dmonte et al., 2024), enabling the creation of many notable benchmarks for these tasks, such as FEVER (Thorne et al., 2018a, b), WikiFactCheck-English (Sathe et al., 2020), VitaminC (Schuster et al., 2021), and WICE (Kamoi et al., 2023). However, whereas these works treat Wikipedia articles as sets of claims or passages to sample from for dataset curation, this work studies Wikipedia articles as whole, structured documents—relied upon as trustworthy sources for information-seeking tasks. [^2] Of these, WICE is most similar to our work. Appendix C has a detailed discussion of differences.
First, we ask to what extent claims in Wikipedia are grounded. Acknowledging Wikipedia’s distinction between an article’s lead section (i.e. the intro) and its body, we are the first to jointly explore both how claims in the lead are grounded in the body (article-internal support) and how claims in the body are in turn grounded in publicly accessible cited sources (article-external support). Second, we ask how effectively standard retrieval methods can recover evidence about these claims—either from the body (for lead claims) or from cited source documents (for body claims). In answering these questions, we make the following contributions:
  • We release PeopleProfiles, a new dataset of structured Wikipedia claim support judgments for all lead claims and all body claims with scrapable citations from 1.5K articles about people, covering nearly 50K lead claims and 100K body claims with fine-grained scalar support labels and associated evidence.
  • We show that: (1) a surprising proportion of lead claims ($\sim 20\%$) are unsupported by the body of the same article; (2) an even higher proportion of body claims ($\sim 27\%$) are unsupported by scrapable sources.; and (3) $>80\%$ lead claims cannot be directly grounded in these sources via supporting body sentences.
  • We show that even in Wikipedia, evidence for these claims is often complex, involving multiple premises, and that retrieval of such evidence remains challenging.
Figure 1: An example of the multi-level structure of PeopleProfiles annotations. Claims in the lead of a Wikipedia article (top left) are supported by sentences in the body (bottom left), whose claims in turn are supported by evidence in cited sources (right). Prior work on Wikipedia claim verification has not considered this structure.
2 Data Collection
Methodology
We obtain evidence for Wikipedia claims and scalar [-1,1] judgments of the degree of support/refutation for those claims given that evidence. [^3] While refutation is unlikely in Wikipedia, we wanted to capture the rare cases where it occurs. We divide annotation into two phases—one for claims appearing in the article’s lead and a second for claims appearing in its body. This is motivated by the different guidelines Wikipedia establishes for these two parts of an article: while citations are required for contentful claims in the body (e.g. quotations, statistics), [^4] https://en.wikipedia.org/wiki/Wikipedia:When_to_cite “it is common for citations to appear in the body and not the lead,” since “significant information should not appear in the lead if it is not covered in the remainder of the article.” [^5] https://en.wikipedia.org/wiki/Wikipedia:Manual_of_Style/Lead_section Thus, for lead claims, we seek evidence in the body, and for body claims, we seek evidence in cited sources. Following prior work (Kamoi et al., 2023), we define the evidence for a claim as a set of (possibly non-contiguous) sentences. We annotate up to 3 sentences that together provide the strongest evidence for or against each target claim.
Claims
We adopt the view advocated in work on claim decomposition that the appropriate units for assessment of evidential support are subclaims, i.e., sub-sentence-level statements expressing an atomic proposition (Kamoi et al., 2023; Min et al., 2023; Wanner et al., 2024a, b; Gunjal and Durrett, 2024, i.a.). [^6] When we refer to claims in this work, we mean subclaims. We use the “DnD” method of Wanner et al. (2024b) to jointly decompose each Wikipedia sentence into two sets of subclaims: a contextualized set decomposed from the sentence alone and a decontextualized set that inserts into each subclaim relevant extra-sentential context (e.g. to resolve pronouns). Following Wanner et al., we use GPT-4o-mini (OpenAI, 2024) to perform the decomposition. [^7] See Appendix A for prompts. Annotators can see both versions of a subclaim when assessing its support.
Train Dev Test
Articles 965 256 264
Lead Claims 30,331 9,272 9,351
Body Claims 60,107 19,712 18,712
Sources 10,539 3,298 3,485
Table 1: PeopleProfiles summary statistics.
Data Source
We select for annotation Wikipedia articles of notable people (entities) studied in prior work on claim verification, including all those from Min et al. (2023) and Jiang et al. (2024), yielding 1,485 entities that represent a range of nationalities and degrees of fame. We obtain the (structured) English articles for each entity from the MegaWika 2.0 dataset (Barham et al., 2023, 2025), including their in-text citations and the citations’ scraped source texts. We annotate claim support for subclaims decomposed from all sentences in articles’ leads and all body sentences that bear citations to publicly accessible sources, as we cannot verify paywalled or print sources at scale—nor can Wikipedia users or RAG-enabled search engines. We use the DeBERTa-based (He et al., 2020) text quality classifier from NVIDIA’s NeMo Curator to filter low-quality sources. [^8] https://github.com/NVIDIA/NeMo-Curator We divide entities into train (65%) and dev and test (35%) splits via stratified sampling on the number of lead subclaims.
Pilot Annotation
To ensure high-quality automatic annotation on the full entity set, we conduct a pilot human annotation on a set of 160 body claims obtained from 10 entities, divided into 3 batches. Each batch was annotated with two-way redundancy by three authors, using an interface and instructions we designed for the task (see Appendix A). We assess inter-annotator agreement on support judgments using Krippendorff’s $\alpha$ (Krippendorff, 2018) and on the selected sets of evidence sentences using average pairwise $\text{F}_{1}$ (with exact sentence match), obtaining $\alpha=54.3$ and $\text{F}_{1}=53.8$. We then use these results to guide prompt engineering for the bulk annotation, assessing GPT-4o-mini on the same examples. Model annotations from our final (best) prompt yield $\alpha=64.6$ and $\text{F}_{1}=54.1$ when included with the original human ones—indicating that GPT-4o-mini achieves inter-human agreement levels.
Bulk Annotation
Using GPT-4o-mini with the same prompt, we collect support and evidence annotations on all 1,485 entities. Table 1 shows statistics of the resulting PeopleProfiles dataset.
Figure 2: Kernel density estimation plots for Wikipedia lead/body claim support in the PeopleProfiles dev split. We find that many claims are not fully grounded.
3 Analysis & Experiments
3.1 Claim Support
A significant fraction of lead and body claims are unsupported.
Figure 2 plots lead and body claim support distributions for the PeopleProfiles dev split. We observe strong bimodality in both distributions, with high density around both full support (1.0) and no support (0.0). Indeed, 19.3% of lead claims are judged unsupported ($\leq 0$) by the body text and 26.5% of body claims by their cited source text(s). Inspection reveals that, contrary to guidelines, many leads make assertions attested nowhere else in the article—notably, about birth and death date and location—while other unsupported claims present inherently difficult attribution problems (e.g. nickname origins). Similarly, many body claims assert propositions unattested in publicly accessible sources: numerous articles extensively cite copyrighted books or paywalled articles which, while clearly legitimate, places hard limits on the amount of content that can be readily verified by (human or machine) readers.
Support does not robustly propagate from sources up to lead claims.
We directly annotate lead claim support given body evidence, but we can further consider how strong the support is for that evidence based on cited sources. We consider two methods of computing a support score for a body evidence sentence given the support scores of its decomposed claims, either taking the scores’ mean or their product (clipping scores $<0$ to 00). We can then compute an overall score for an evidence set via the same aggregations applied to the set, yielding 4 possible overall scores. Broadly we find that (1) most lead claims (82%) do not ground out in source evidence because their body evidence sentences lack citations; (2) of those that do, average overall evidence scores are very modest (e.g. 0.41 when using mean-mean aggregation). [^9] Appendix B plots these overall score distributions.
Task Model N@5 R@5 R@10
LEAD ColBERTv2 52.59 57.90 68.18
Stella-1.5B-v5 30.03 38.03 51.35
BM25 49.92 56.02 66.01
+Rerank Rank1 60.55 63.25 66.01
BODY ColBERTv2 70.02 76.37 87.16
Stella-1.5B-v5 49.37 60.89 78.15
BM25 61.70 68.21 80.24
+Rerank Rank1 73.02 76.18 80.24
Task Model N@5 N@100 R@100
ENTITY ColBERTv2 24.53 39.65 62.52
Stella-1.5B-v5 13.56 28.40 57.96
BM25 14.59 30.08 56.67
+Rerank Rank1 20.54 34.28 56.67
Table 2: Evidence retrieval results for LEAD, BODY, and ENTITY. Best first-stage results are bolded. “+Rerank” is reranked BM25 results ($k=10$ for LEAD and BODY; $k=100$ for ENTITY). N=NDCG; R=Recall.
3.2 Evidence Retrieval
We consider three evidence retrieval tasks:
  1. 1.
    LEAD: Retrieve the body evidence sentences for a given lead claim
  2. 2.
    BODY: Retrieve the evidence sentences from a single cited source for a given body claim
  3. 3.
    ENTITY: Retrieve all evidence sentences annotated in all cited sources for a given entity
[…]
For LEAD and BODY, we report recall@{5,10} and NDCG@5 results on the PeopleProfiles test set using several widely used retrieval models: BM25 (Robertson et al., 1995), ColBERTv2 (Khattab and Zaharia, 2020; Santhanam et al., 2022), and Stella-v5 1.5B (Zhang et al., 2024). Since ENTITY requires retrieval of many more sentences, we instead report recall@100, as well as NDCG@{5,100}.
Main Results
Table 2 reports the main results for all three models on all three tasks. We consistently obtain our best results with ColBERTv2, which shows 2+ point gains across metrics on LEAD and ENTITY, and 6+ point gains on BODY relative to the next best method.
Evidence retrieval difficulty increases with query scope.
We observe wide variability in the difficulty of different tasks, with highest NDCG@5 scores on BODY, followed by LEAD and then by ENTITY. Intriguingly, this ranking tracks the granularity of claims/queries, where body claims (BODY) tend to provide the most detailed information, lead claims (LEAD) present key high-level facts, and entity-level queries (ENTITY) represent a limiting case—seeking generic biographical information. Intuitively, highly specific body claims likely bear greater lexical and semantic similarity to their supporting material than the higher-level claims of leads or the entity-level queries do to theirs.
Task #Sents N@5 R@5 R@10
LEAD 1 88.20 88.90 72.39
2 57.29 68.42 52.77
3 33.72 46.83 32.13
BODY 1 85.71 91.93 75.78
2 66.13 79.68 59.71
3 48.28 66.01 45.98
Table 3: Retrieval results for ColBERTv2 broken down by number of evidence sentences. Retrieval performance drops sharply as # evidence sentences increases.
Evidence retrieval difficulty increases with evidence complexity.
Table 3 presents retrieval results on LEAD and BODY broken down by number of annotated evidence sentences. Whereas retrieval performance is strong for single-sentence evidence, we observe double-digit drops in moving to 2- and 3-sentence evidence sets. This may be explained by the fact that evidential support is often compositional, requiring integration of independently non- or weakly supporting pieces of evidence via inference rules. Addressing this by indexing larger passages, though tempting, would severely curtail the ability to localize the evidence: the average distance between evidence sentences for dev set body claims with multi-sentence evidence sets is 8.7 sentences, expanding to 14.6 for lead claims. That we find evidence dispersed across multiple sentences even in Wikipedia indicates that the need to handle compositional evidence is not a niche concern.
Reasoning rerankers help.
The above observations suggest that effective retrieval of complex evidence demands more sophisticated methods than lexical or semantic similarity match. Recent work shows that reasoning-based rerankers achieve substantial gains on other complex retrieval tasks (Weller et al., 2025; Shao et al., 2025; Zhuang et al., 2025). Accordingly, we leverage Rank1-7B, a pointwise reranker based on Qwen 2.5 7B (Qwen et al., 2025) distilled from 635K reasoning traces for MS MARCO relevance judgments produced by DeepSeek R1 (Guo et al., 2025). We use Rank1 to rerank the top 10 evidence sentences from BM25 for LEAD and BODY and the top 100 for ENTITY. Results are in Table 2’s “+Rerank” rows, where we find large gains over first-stage retrieval across all metrics—pointing to a vital role for reasoning-based rerankers in complex evidence retrieval. Table 4 has fine-grained results.
4 Conclusion
We have presented a study of evidential support in Wikipedia and have introduced PeopleProfiles, a large new resource of fine-grained, multi-level support annotations on 1,500 Wikipedia articles and their cited sources. We have shown that: (1) a sizable fraction of Wikipedia lead claims are unsupported by their body sections, and Wikipedia body claims by their (publicly accessible) cited sources; (2) evidence retrieval for these claims grows much more challenging as query scope and evidence complexity increases; and (3) new reasoning-based rerankers enable much more effective retrieval of complex evidence. We release PeopleProfiles to aid future work on claim verification and on furthering understanding of Wikipedia as a knowledge base for modern NLP.
Limitations
We acknowledge several limitations of our work. First, PeopleProfiles focuses only on Wikipedia articles about people. We chose this focus because biographies tend to have a higher proportion of straightforward, uncontroversial facts relative to other domains (e.g. concepts or events). However, it is possible the support distributions or the difficulty of evidence retrieval for articles in these other domains could differ from what we observe here. Second, as we emphasize in the main text, our claims about evidential support extend only to publicly accessible, digital sources—those that a human or machine reader could readily use to verify an article’s claims. We therefore cannot make conclusions about support across all source types in Wikipedia. Third, we leverage GPT-4o-mini as an annotator to facilitate our large-scale bulk data collection. While the agreement we observe between this model and our human annotations is strong (§2), LLMs have their own response biases and may not be fully calibrated when providing scalar judgments (Lovering et al., 2024).
Finally, a major challenge for all NLP work on citation is that real-world citation practices routinely deviate from the highly idealized model, widely adopted by NLP practitioners, in which each citation-worthy sentence is expected to bear a citation. In practice, a citation may be given (e.g.) only once at the beginning of a paragraph if it is clear that the remainder of the paragraph concerns the same source. Alternatively, several citations may be provided together only in the final sentence of a paragraph. Such practices may obscure the mappings between claims and their source material, but they are the norm and it is incumbent on the NLP community to find better ways to accommodate them.
Since we restrict ourselves to body sentences that do bear citations, we cannot draw conclusions about the (likely substantial) number of sentences whose source may appear as a citation on some other sentence. However, we suspect our estimate of the proportion of supported body claims decomposed from citation-bearing sentences is an upper bound on the proportion we would obtain for all citation-worthy sentences.
Ethics
PeopleProfiles’s use of sources from MegaWika and our release of this data (via a CC-BY-4.0-SA license) is consistent with MegaWika’s own CC-BY-4.0-SA license. Our principle transformation of the original Wikipedia articles consists in the decomposition of claims, which is performed by an LLM (GPT-4o-mini), and which can result in subclaims that misrepresent the article’s original content and thus (potentially) facts about the subject. Although our claim decompositions are highly faithful to the original texts, users of PeopleProfiles should be aware of this possibility.
Acknowledgments
The authors would like to thank Alex Martin for help testing out the annotation interface, as well as colleagues at the Human Language Technology Center of Excellence for general comments and feedback on this work.
Appendix A Data Collection
A.1 Annotator Demographics
The first three authors of this work, all native English-speaking graduate or professional NLP researchers, conducted the human pilot annotations. These authors also jointly produced the annotation instructions beforehand. None was compensated beyond their co-authorship on this work.
A.2 Claim Decomposition
Decomposition is the process of breaking down sentences into simpler, atomic components—aiming to isolate individual, independent claims for downstream applications. A common approach to claim decomposition uses LLMs to segment a sentence into independent facts, each containing one “piece” of information. However, these subclaims can be ambiguous, with vague references that are uninterpretable without the context of the document. The process of decontextualization mitigates this issue by rephrasing a subclaim such that it is fully intelligible as a standalone statement, without the original document as context. These two processes are complementary: decomposition divides sentences into smaller parts, whereas decontextualization adds information.
We use the “DnD” decomposition and decontextualization method introduced by Wanner et al. (2024b), which uses an LLM prompt-based method for obtaining subclaim decompositions and the corresponding decontextualized subclaims. We decompose and decontextualize sentences from the original Wikipedia page, either from the lead (in the LEAD task) or body (in the BODY task), and provide the lead paragraph (LEAD) or additionally the body paragraph from which the claim originates (BODY) as context for decontextualization. During the pilot annotation, annotators are able to toggle between the subclaim and its decontextualized version to select evidence sentences that together provide the strongest evidence (either supporting or refuting) for the subclaim. Finally, after identifying evidence, annotators determine a support score for the subclaim given that evidence. The bulk annotation provides only the decontextualized subclaim as lead or body claim. Following Wanner et al., we use GPT-4o-mini (OpenAI, 2024) to perform DnD.
A.3 Annotation Interface
The annotation interface used for the human annotation is shown in Figure 3. The full, sentence-split text of a cited source article is shown on the far left. All of the subclaims decomposed from a single Wikipedia body sentence citing that source article are shown in a vertical list of tiles on the far right, with the currently selected subclaim displayed in the top middle part of the screen (to the right of “Claim:”). Here, annotators can toggle between the original and decontextualized versions of the subclaim using the D toggle shown above the subclaim, with differences (additions, deletions) between the decontextualized and original versions shown in blue and red. Annotators can also display the sentence that the current subclaim was decomposed from, along with its full Wikipedia context, by clicking the More Info toggle in the top right.
Several checkboxes are also shown above the subclaim to enable annotators to indicate that:
  • The source text is uninterpretable or otherwise low quality (Bad Source)
  • The subclaim is unfaithful to the meaning of the original sentence (Bad Decontextualization)
  • It is simply too difficult to determine how the current subclaim relates to the source material—e.g. because the source document is too technical for the annotator to understand (I’m Uncertain)
Annotators select up to three sentences from the source text on the left that together provide the strongest evidence (either supporting or refuting) for the target subclaim. We chose a maximum of three sentences because this enabled adequate coverage of the evidence for the vast majority of claims while keeping the task tractable for annotators.
Finally, the blue box (bottom middle) is used to specify the support score for the currently selected subclaim, given the identified evidence. After selecting evidence and providing a support score for all subclaims (toggling between them using the NEXT and BACK buttons on bottom), annotators submit their work via the SUBMIT button.
Figure 3: Annotation interface for the human pilot annotation. Detailed description can be found in Appendix A.2.
A.4 Prompts and Hyperparameters
The prompt used for bulk annotation with GPT-4o-mini is shown in Figure 5 through Figure 9 (divided over multiple pages due to the length of the instructions). This prompt was selected based on highest agreement with the human pilot annotations after numerous manual iterations on other prompts. We used gpt-4o-mini-2024-07-18, the most recent version of the model available. Annotations were generated with temperature 0, with a limit of 2K output tokens to accommodate source texts of up to 126K tokens. Source texts exceeding this limit were truncated, though this was rarely required.
Appendix B Experimental Details and Additional Results
B.1 Qrels for ENTITY
For the ENTITY task in §3, we assign fine-grained relevance labels to sentences in the source documents for a given entity based on (1) how strongly they support a Wikipedia body claim, (2) how many body claims they support, (3) how strongly they support lead claims via body claims, and (4) how many lead claims they support.
Given an article for entity $E$, a sentence $S_{B}$ in the article’s body, a sentence $S_{S}$ in some cited source, and a claim $C$, we define the following:
  • $lead_{E}(S_{B})$: the set of lead claims that have $S_{B}$ in their (body) evidence set
  • $body_{E}(S_{S})$: the set of body claims that have $S_{S}$ in their (source) evidence set
  • $support(C)$: the support score for a claim $C$
  • $sent(C)$: the sentence that claim $C$ was decomposed from
Letting $C_{B}$ be a body claim and $C_{L}$ be a lead claim, we then define the relevance of a source sentence $S_{S}$ to a query $Q_{E}$ about entity $E$ as the following weighted sum:
$Rel(Q_{E},S_{S})$ $=\sum_{C_{B}\in body_{E}(S_{S})}w_{C_{B}}$
$\cdot\text{abs}(support(C_{B}))$
$w_{C_{B}}$ $=1+\sum_{C_{L}\in lead_{E}(sent(C_{B}))}$
$\text{abs}(support(C_{L}))$
Intuitively, $Rel(Q_{E},S_{S})$ is a weighted sum of the absolute values of the support scores of all body claims ($C_{B}$’s) that $S$ is evidence for ($body_{E}(S_{S})$). We use the absolute value of the support score because $S$ is equally important as evidence regardless of whether it is supporting or refuting evidence.
The weight $w_{C_{B}}$ associated with each body claim $C_{B}$ is 1, plus the sum of (absolute values of) support scores of all lead claims for which $sent(C_{B})$—the sentence $C_{B}$ was decomposed from—provides evidence. This rewards $S$ for indirectly supporting a lead claim $C_{L}$ via a body claim ($C_{B}$), proportional to the degree of support for $C_{L}$. The motivation here is simply that (1) lead claims typically represent more important facts about an entity than body claims, and thus sentences that (indirectly) provide evidence for them should be rewarded, and (2) that reward should be proportional to the degree of support.
We note that this is a somewhat heuristic weighting scheme, as $C_{B}$ is given credit merely for being decomposed from a sentence that supports a lead claim $C_{L}$—even if a different claim ($C_{B}^{\prime}$) decomposed from the same sentence provides the bulk of the evidence for $C_{L}$. Collecting further annotations to enable more precise assignment of relevance scores is a direction we are pursuing for future work.
B.2 Fine-Grained Reranking Results
Table 4 shows the BM25 and Rank1 results from Table 2 broken down by number of evidence sentences in the gold annotations for each query (note: R@10 results are omitted, as they are unchanged by reranking the top-10 sentences). These results convincingly demonstrate that the gains brought by leveraging a reasoning model (Rank1) for reranking are not limited to the “easy” cases of single-sentence evidence sets but robustly extend to multi-sentence evidence sets as well.
Task #Sents Model NDCG@5 R@5
LEAD 1 BM25 76.14 86.64
Rank1 84.49 90.67
2 BM25 53.11 47.28
Rank1 61.32 63.47
3 BM25 25.34 27.40
Rank1 35.40 35.28
BODY 1 BM25 75.78 85.71
Rank1 85.72 90.71
2 BM25 59.71 66.13
Rank1 71.90 75.38
3 BM25 45.98 48.28
Rank1 58.10 58.60
Table 4: Gains from reranking the top-10 BM25 evidence sentences for LEAD and BODY using Rank1, broken down by number of evidence sentences for the query/claim. Rank1 shows large gains in all cases.
B.3 Evidence Propagation
§3 briefly presents some analysis on support for the body evidence for a given lead claim. There, we say that we compute an evidence score for a body sentence by taking either the mean or the product of the support scores for its constituent claims (clipping scores $<0$ to 0 in the latter case). We can then compute an overall evidence score for an evidence set by taking the mean or product of the per-sentence scores. Figure 4 plots distributions of overall evidence scores in the PeopleProfiles dev split when applying both mean (blue) and product (orange) aggregation over claims and then (in both cases) applying mean aggregation over sentences. In both cases, we find obtain very middling overall evidence scores—an average of 0.41 for mean and an average of just 0.12 for product.
Figure 4: Distribution of overall evidence scores for PeopleProfiles dev split body evidence with mean- (blue) and product-based (orange) aggregation of body claim support scores for each evidence sentence. See Appendix B.3.
B.4 Retrieval Model Details
For BM25 (no parameters), we use the implementation provided in the bm25s library (Lù, 2024) with default settings. We access Stella-1.5B-v5 (1.5 billion parameters) through the sentence-transformers library with default settings (i.e. no hyperparameter search was performed). Finally we access ColBERTv2 (jinaai/jina-colbert-v2 on HuggingFace; 559M parameters) via the ragatouille library [^10] https://github.com/AnswerDotAI/RAGatouille, leveraging FAISS for indexing (Johnson et al., 2019), and again using default settings. Neither Stella-1.5B-v5 nor ColBERTv2 were fine-tuned on PeopleProfiles. All experiments were carried out on a single NVIDIA A100 GPU except the reranking experiments, for which four A100s were used. All main text results reflect single runs.
The prompts for reranking evidence with Rank1-7B are provided in Figure 10 and Figure 11. Outputs were generated with temperature 0. Context size was set to 16K tokens, with a maximum of 8192 output tokens.
B.5 Use of AI Assistants
No AI assistance was used in the ideation or in the writing of this paper. GitHub Copilot was used to assist in writing the code for some of the experiments and analysis.
Appendix C Further Discussion of Related Work
In §1, we note that the resource most similar to PeopleProfiles is the WiCE dataset from Kamoi et al. (2023), a textual entailment dataset using text-citation pairs from Wikipedia. Here, we discuss some of the key differences between our PeopleProfiles and WiCE, summarized in Table 5. First, support scores in PeopleProfiles are scalar, rather than categorical (SUPPORTED, PARTIALLY-SUPPORTED, NOT-SUPPORTED), as in WiCE, which enables finer-grained analysis of partial support (see §3). Furthermore, PeopleProfiles includes article-internal annotations of claim support (LEAD) in addition to article-external annotations (BODY), whereas WiCE contains only the latter. To our knowledge, ours is the first work to have both types of claim support annotations. We also annotate all lead sentences and all body sentences with attached citations, with WiCE opting to annotate only the SIDE subset (Petroni et al., 2022), containing citations unlikely to support the claim. Although PeopleProfiles annotations are automated by an LLM rather than curated by humans, this allows us to have a dataset over 20×20$\times$20 × as large as WiCE.
Dataset Characteristic Split WiCE PeopleProfiles (Ours)
Support Scores Categorical Scalar
Article-internal grounding annotations
Article-external grounding annotations
Subset of article-external subclaims annotated SIDE subset (Petroni et al., 2022) All available
Annotations per subclaim Train 3 human 1 LLM
Dev 5 human 1 LLM
Test 5 human 1 LLM
Number of body subclaims Train 3,470 60,107
Dev 949 19,712
Test 958 18,712
Table 5: Comparison of dataset characteristics between WiCE and our proposed PeopleProfiles.
PeopleProfiles Annotation Prompt In this task, you will be shown a claim along with a list of sentences representing a document that might provide evidence for the claim. Given this information, you will perform two steps, described below. For both steps, rely on the following two definitions of evidence: Definition 1: “Supporting evidence”: A set of sentences S provides supporting evidence for a claim c if, supposing the contents of S were true, it would give you greater reason to believe that c is true, all else equal. Definition 2: “Refuting evidence”: A set of sentences S provides refuting evidence for a claim c if, supposing the contents of S were true, it would give you greater reason to believe that c is false, all else equal. Step 1: Select 0, 1, 2, or *at maximum* 3 sentence(s) from the document that provide the strongest supporting evidence or refuting evidence for the claim. If no sentences in the document provide evidence, do not select any sentences. Additional guidelines for Step 1: (a) You may need to use logic and common sense to *infer* that a sentence provides evidence for the claim. For example, you can use common sense to assume that a person wearing reading glasses struggles with their sight. (b) Do not assume any parts of the claim are common knowledge. You must find evidence for all parts of the claim. For example, if the claim states that Vidya, the English chef, has poor vision, you would need to find evidence that Vidya is English and a chef, as well. (c) A sentence might provide evidence for the claim only when combined with other sentences. For example, if Sentence A states Bob is married to Mary, and Sentence B states that Mary is a doctor, Sentences A and B together provide supporting evidence for the claim that Bob has a doctor in his family. (d) Please make sure the entities and events in your selected sentences match those in the claim. For example, dates and names, as determined by the rest of the document, should match the claim; else, the sentences do not provide evidence.
Figure 5:
PeopleProfiles Annotation Prompt, continued Step 2: Given your selected set of sentences from Step 1, score the degree to which these sentences (taken together) support or refute the claim. Determine the score according to the following definition of a scale from -1 to 1: -1: The claim is *fully refuted*: The claim would have to be false, supposing the sentences you selected were true. Scores between -1 and 0 (-0.9, -0.8, -0.7, -0.6, -0.5, -0.4, -0.3, -0.2, -0.1): The claim is *partially refuted*. The claim would have to be false, but some parts are likely true. 0: The claim is neither supported nor refuted. It is equally likely to be true or false. Scores between 0 and 1 (0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9): The claim is *partially supported*. The claim is likely partially true, with missing evidence. No parts of the claim are likely to be false. 1: The claim is *fully supported*: The claim would have to be fully true, supposing the sentences you selected were true. Additional guidelines for Step 2: (a) Use only the content of your selected sentences to make your judgment. Do not use any knowledge you may already have about the claim, nor any context from other sentences in the document. For example, even if you know that London is in England, or it is stated elsewhere in the document, you cannot judge that detail of the claim as supported unless it is stated in your selected sentences. (b) As in Part 1, do not assume any parts of the claim are common knowledge. Assign the score based on all parts of the claim, even if they seem obviously true or false. (c) The document might only contain evidence for a similar but distinct claim. For example, if the strongest evidence states that the president ate at a restaurant on a Friday, this is not refuting evidence for the claim that the president ate at a restaurant on Tuesday; in fact, there is no evidence to support or refute the claim.
Figure 6:
PeopleProfiles Annotation Prompt, continued Below are 10 examples of scoring sentences that have already been selected from a document as supporting or refuting evidence for a claim: ###Example 1### Claim: "Methane Momma is a short film directed by Alain Rimbert." Selected sentences: ["Well, good news 2̆013 last week, in the middle of one of the worst heat waves that New York has seen in recent memory, a pajama-clad (and still ripped) Van Peebles entered ex-Sun Ra bandmember Spaceman’s Harlem-based studio and recorded his last takes on the rambling poem he’s entitled M̈ethane Momma."̈] Score: -0.7 ###Example 2### Claim: "Raj Kapoor was hospitalised for about a month." Selected sentences: ["Suddenly, Kapoor collapsed, and was rushed to the All India Institute of Medical Sciences for treatment.", "The country’s top cardiologists tried their best, but could not save him."] Score: -0.1 ###Example 3### Claim: "Ottawa is a city located in the province of Ontario, Canada, and is where Matthew Perry attended school." Selected sentences: [] Score: 0 ###Example 4### Claim: "Paul Thomas Anderson registered himself with the Writers Guild of America under the name ’Paul Anderson.’" Selected sentences: [] Score: 0 ###Example 5### Claim: "There were exile forces opposing Idi Amin’s regime." Selected sentences: ["Since leading his guerrilla forces to Kampala in 1986, his most impressive flexibility has been his capacity to present two concurrent faces: one is that of the democratic reformer, the other is of the fearsome military ruler.", "The former is the saviour of Uganda’s post-colonial collapse under presidents Milton Obote and Idi Amin, patron of democracy, and emancipator of woman and ethnic and religious minorities."] Score: 0.1
Figure 7:
PeopleProfiles Annotation Prompt, continued ###Example 6### Claim: "Margaret Rose Vendryes wrote about Richmond Barth0̆0e9’s work further in her 2008 book." Selected sentences: ["By coincidence, Dr. Vendryes was the Schomburg’s scholar-in-residence and was researching her Princeton doctorate thesis on Barthe, which evolved into her landmark book Casting Feral Benga: A Biography of Richmond Barth0̆0e9’s Signature Work."] Score: 0.3 ###Example 7### Claim: "Margaret Rose Vendryes gave a lecture in 2015." Selected sentences: ["This Thursday, February 5 at the Jepson Center, Dr. Vendryes will give the opening lecture for the exhibition."] Score: 0.5 ###Example 8### Claim: "The exhibit presented by The New York Public Library for the Performing Arts was extensive." Selected sentences: ["Curated by Doug Reside, the Lewis B. and Dorothy Cullman curator of the library’s Billy Rose Theatre Division, the installation will run through March 31, 2020, and feature original costumes, set models, and archival video tied to Prince’s productions, including models for several productions.", "The full display will honor the more than six-decade legacy of Prince.", "An open cabaret stage will allow viewers to perform songs from his shows or record their own stories about their experience with Prince’s theatrical work to add to the live nature of the homage."] Score: 0.7 ###Example 9### Claim: "The location of Matthew Perry’s funeral was Forest Lawn Memorial Park (Hollywood Hills), a cemetery." Selected sentences: ["Photo: David M. Benett/Dave Benett/Getty Matthew Perry’s loved ones gathered for the actor’s funeral on Friday.", "The service was held at Forest Lawn Memorial Park in Los Angeles near Warner Bros. Studios.,"] Score: 0.9 ###Example 10### Claim: "The promotional video was 60 minutes long." Selected sentences: ["Microsoft made a c̈yber sitcomẗo promote it.", "The final product [debuted on VHS on August 1, 1995](https://books.google.com/books?id=0QsEAAAAMBAJ& lpg=RA1-PA62&dq=matthew%20perry%20jennifer%20aniston%20windows%2095&pg=RA1-PA62#v=onepage&q&f=false), satisfying everybody who wished Friends were an hour long, had four fewer friends, and involved a guide to file management."] Score: 1
Figure 8:
PeopleProfiles Annotation Prompt, continued Finally, here are the claim and list of document sentences for your task: Claim: <subclaim> Document sentences: <numbered source sentences> Write your response in a dictionary in the format shown below. Write the dictionary and nothing else. Dictionary format: "sentences": [ "[<sentence number>] <sentence selected from document>", …, ], "score": <number between -1 and 1> ###Your Task### Selected sentences and score in dictionary form:
Figure 9:
PeopleProfiles Evidence Reranking Prompt: BODY and LEAD The following is a claim: <claim> A relevant passage provides supporting or refuting evidence for the claim.
Figure 10:
PeopleProfiles Evidence Reranking Prompt: ENTITY I am writing an encyclopedia article about the following person: <entity>. A relevant passage contains noteworthy biographical facts about this person. For example, a passage containing facts about this person’s early life, education, career, or death is relevant.
Figure 11: